OpenStudy (anonymous):

Triangle ABC is similar to triangle DEF and has a ratio of similarity of 5:2. If side AB measures 30 mm, how long is side DE?

OpenStudy (callisto):

Triangle ABC is similar to triangle DEF and has a ratio of similarity of 5:2, \[\frac{5}{2} = \frac{30}{DE}\]Can you solve DE?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

can i ask you one more please

OpenStudy (anonymous):

In a 30°- 60°- 90° right triangle, the shorter leg is 4 cm. How long are the other two sides of the triangle?

OpenStudy (callisto):

 sin30 = 4/ a a = 4/ sin30 = ? tan30 = 4/b b= 4/tan30 = ?